# Convergence of Decentralized Stochastic Subgradient-based Methods for Nonsmooth Nonconvex functions

### v3

@@ 5 Numerical Experiments
Figure 1: Numerical performance comparison of DSGD, DSGD-M, and DSignSGD in training ResNet50 on CIFAR-100 dataset.
@@
Figure 2: Numerical performance comparison of DSGD, DSGD-M and DSignSGD in training ResNet50 on CIFAR-10 dataset.
@@ 5.1 Experiment settings
We conduct a series of image classification tasks with CIFAR-10 and CIFAR-100 datasets [67], which both consist of $50000$ training samples and $10000$ test samples. ResNet-50 [68] is employed as our nonsmooth neural network. The decentralized network topology is set as ring-structure with $8$ agents, and the mixing matrix is chosen as the Metropolis weight matrix [54]. In each experiment, we divide the training dataset into 8 equal parts, with each part serving as the local dataset for each agent. During any epoch of training, the local dataset of each agent itself remains unchanged, while the mini-batch is selected by random reshuffling to ensure compliance with the sampling scheme in Assumption 4.2.
@@ 5.2 Performance comparison between DSGD, DSGD-M, and DSignSGD
We present the numerical results of the above-mentioned decentralized stochastic sub-gradient-based methods in Figure 1 and 2, including the test accuracy, train loss, and consensus error. It is worth mentioning that the consensus error at the $k$-th iteration is measured by $\frac{1}{\sqrt{d}}\|{\bm{X}}_{k}({\bm{I}}_{d}-{\bm{P}})\|_{\mathrm{F}}$. As illustrated in Figure 1 and 2, all the compared decentralized stochastic subgradient-based methods are able to train the ResNet-50 network to a high accuracy. Moreover, we can conclude from Figure 1(a) and 2(a) that DSignSGD achieves similar test accuracy as DSGD and better test accuracy compared to DSGD-M, while showing superior performance to DSGD and DSGD-M in the aspect of train loss from Figure 1(b) and 2(b). Furthermore, in all the compared methods, the consensus error diminishes to 00 throughout the iterations. These results illustrate that our framework is able to yield efficient stochastic subgradient-based methods for decentralized training tasks.
@@ 5.3 Performance comparison between DSGD and DSGD-T
The numerical results on the CIFAR-100 and CIFAR-10 datasets are presented in Figure 3 and 4, respectively. It can be concluded that DSGD-T demonstrates slightly better performance than DSGD in all aspects, especially in train loss. Therefore, we can conclude that DSGD-T achieves high efficiency in decentralized training tasks while enjoying convergence guarantees from our framework (DSM).
@@
Figure 3: Numerical performance comparison of DSGD and DSGD-T in training ResNet50 on CIFAR-100 dataset.
@@
Figure 4: Numerical performance comparison of DSGD and DSGD-T in training ResNet50 on CIFAR-10 dataset.
@@ 6 Conclusion
Our work sets several directions for future research in the field of nonsmooth nonconvex decentralized optimization. For example, non-asymptotic convergence analyses have been established in several existing works [28, 30], but their proof techniques cannot be extended to non-Clarke-regular cases. How to establish the non-asymptotic convergence rate for (DSM) is worthy of investigating. Further attention is also needed on the convergence of decentralized methods with time-varying networks, as well as those that incorporate asynchronous or compressed communication. Additionally, we are also eager to explore routes that presume relaxed conditions for convergence analysis of decentralized methods in solving (DOP). For instance, closed-measure approaches [60, 44, 42] are developed for establishing the convergence properties of stochastic subgradient-based methods, with relaxed conditions on the step-sizes and evaluation noises. How to employ these closed-measure approaches for analyzing decentralized stochastic subgradient-based methods is worthy of investigating.
@@ Acknowledgement
$\mathcal{U}_{k}({\bm{x}})$ $:=\left(D_{F_{1},\mathcal{B}_{1,k+1}}({\bm{x}}_{(1)}),\ldots,D_{F_{d},\mathcal{B}_{d,k+1}}({\bm{x}}_{(d)})\right)^{\top}$
Note that $(F_{i}(\cdot,{\bm{s}}_{l}),D_{F_{i}}(\cdot,{\bm{s}}_{l}))$ has a $\mathcal{C}^{r}$ variational stratification [36, Theorem 4], hence the mappings $\{\mathcal{U}_{k}\}$ are almost everywhere $\mathcal{C}^{1}$ over $\mathbb{R}^{nd}$, denoted by $\mathcal{U}_{k}({\bm{x}})=\{u_{k}({\bm{x}})\},$ for a.e. ${\bm{x}}\in\mathbb{R}^{nd}$. Furthermore, let
$\mathcal{Y}$ :=$:=\left\{{\bm{x}}\in\mathbb{R}^{nd}:D_{F_{i}}({\bm{x}};{\bm{s}}_{l})\neq\{\nabla F_{i}({\bm{x}},{\bm{s}}_{l})\},\text{ for some }i\in[N]\text{ and some }{\bm{s}}_{l}\in\mathcal{S}_{i}\right\},$
@@
is an open covering of $\{{\bm{x}}\in\Gamma_{k}:\det({\bm{J}}_{c,k}({\bm{x}}))\neq 0\}$. Based on Lindelof’s lemma [69], there exists countable $\{{\bm{x}}^{i}\}_{i\in\mathbb{N}_{+}}\subseteq\{{\bm{x}}\in\Gamma_{k}:\det({\bm{J}}_{c,k}({\bm{x}}))\neq 0\}$, such that
@@
Let $\mathcal{S}:=\cap_{k\geq p}\mathcal{S}_{k}$, $A_{0}:=A=\mathcal{Y},A_{k+1}:=\{{\bm{x}}\in\mathbb{R}^{nd}:{\bm{P}}{\bm{z}}-cq_{k}({\bm{z}})\in A\}$, and $\hat{A}:=\cup_{k\geq 0}A_{k}$. It can be inferred that $\mathcal{S}$ and $\hat{A}^{c}$ are full-measure subsets in $\mathbb{R}$ and $\mathbb{R}^{nd}$, respectively. For any $c\in\mathcal{S}$, and any ${\bm{x}}\in\hat{A}^{c}$, we have ${\bm{P}}{\bm{x}}-cq_{k}({\bm{x}})\in A^{c}$. Employing Fubini’s theorem, the set $\{c\in\mathbb{R},{\bm{x}}\in\mathbb{R}^{nd},\cup_{k\geq 0}\{{\bm{P}}{\bm{x}}-cq_{k}({\bm{x}})\}\subseteq A^{c}\}$ is full-measure in ×$\mathbb{R}\times\mathbb{R}^{nd}$, and thus, there exists a full-measure subset $\hat{\mathcal{S}}\times\hat{\mathcal{K}}$, whenever $(c,{\bm{x}}_{0})\in\hat{\mathcal{S}}\times\hat{\mathcal{K}}$, it follows that $\cup_{k\geq 0}\{{\bm{P}}{\bm{x}}-cq_{k}({\bm{x}})\}\subseteq A^{c}$, i.e. $\{{\bm{x}}_{k}\}\in\mathcal{Y}^{c}$.
@@
[…]
@@
$\mathcal{U}_{1,k}({\bm{x}},{\bm{y}})$ $:=\left(\partial\hat{\phi}({\bm{y}}),{\bm{0}}\right)$
$\mathcal{U}_{2,k}({\bm{x}},{\bm{y}})$ $:=\left({\bm{0}},\tau({\bm{W}}\otimes{\bm{I}}_{n}){\bm{y}}-\tau\left(D_{F_{1},\mathcal{B}_{1,k+1}}({\bm{x}}_{(1)}),\ldots,D_{F_{d},\mathcal{B}_{d,k+1}}({\bm{x}}_{(d)})\right)^{\top}\right)$
where $\hat{\phi}({\bm{y}}):=(\phi({\bm{y}}_{(1)}),\ldots,\phi({\bm{y}}_{(d)}))^{\top}$, $\phi:\mathbb{R}^{n}\to\mathbb{R}$ is convex and admits a unique minimizer at ${\bm{0}}$. Notice that $(F_{i}(\cdot,{\bm{s}}_{l}),D_{F_{i}(\cdot,{\bm{s}}_{l})})$ has a $\mathcal{C}^{r}$ variational stratification [36, Theorem 4], and $\hat{\phi}$ is differentiable on a full-measure subset of $\mathbb{R}^{nd}$, thus the mappings $\{\mathcal{U}_{j,k}\},j=1,2$ are almost everywhere $\mathcal{C}^{1}$ over $\mathbb{R}^{2nd}$, the corresponding almost-everywhere-$\mathcal{C}^{1}$ mappings are denoted by $u_{j,k}({\bm{z}}),j=1,2$. Furthermore, let
$\mathcal{Y}$ :=$:=\left\{({\bm{x}},{\bm{y}})\in\mathbb{R}^{2nd}:\hat{\phi}({\bm{y}})\text{ is not differentiable },D_{F_{i,l}}({\bm{x}})\neq\{\nabla F_{i}({\bm{x}},{\bm{s}}_{l})\},\text{ for some }i\in[N]\right.$
[…]

### v4

@@ 5 Numerical Experiments
Figure 1: Numerical performance comparison of DSGD, DSGD-M, and DSignSGD in training ResNet50 on CIFAR-100 dataset using random reshuffling strategy.
@@
Figure 2: Numerical performance comparison of DSGD, DSGD-M and DSignSGD in training ResNet50 on CIFAR-10 dataset using random reshuffling strategy.
(a) Test Accuracy
(b) Train Loss
(c) Consensus Error
Figure 3: Numerical performance comparison of DSGD, DSGD-M, and DSignSGD in training ResNet50 on CIFAR-100 dataset using with-replacement sampling strategy.
(a) Test Accuracy
(b) Train Loss
(c) Consensus Error
Figure 4: Numerical performance comparison of DSGD, DSGD-M, and DSignSGD in training ResNet50 on CIFAR-10 dataset using with-replacement sampling strategy.
@@ 5.1 Experiment settings
We conduct a series of image classification tasks with CIFAR-10 and CIFAR-100 datasets [krizhevsky2009learning], which both consist of $50000$ training samples and $10000$ test samples. ResNet-50 [he2016deep] is employed as our nonsmooth neural network. The decentralized network topology is set as ring-structure with $8$ agents, and the mixing matrix is chosen as the Metropolis weight matrix [xiao2006distributed]. In each experiment, we divide the training dataset into 8 equal parts, with each part serving as the local dataset for each agent. During any epoch of training, the local dataset of each agent itself remains unchanged, while the mini-batch is selected by random reshuffling or with-replacement sampling to ensure compliance with the sampling scheme in Assumption 4.2 or 4.3.
@@ 5.2 Performance comparison between DSGD, DSGD-M, and DSignSGD
We present the numerical results of the above-mentioned decentralized stochastic sub-gradient-based methods in Figures 1- 4, including the test accuracy, train loss, and consensus error under two different sampling strategies. It is worth mentioning that the consensus error at the $k$-th iteration is measured by $\frac{1}{\sqrt{d}}\|{\bm{X}}_{k}({\bm{I}}_{d}-{\bm{P}})\|_{\mathrm{F}}$. As illustrated in Figures 1- 4, all the compared decentralized stochastic subgradient-based methods are able to train the ResNet-50 network to a high accuracy. Moreover, we can conclude from Figure 1(a) and 2(a) that when we employ the random reshuffling strategy, DSignSGD achieves similar test accuracy as DSGD and better test accuracy compared to DSGD-M, while showing superior performance to DSGD and DSGD-M in the aspect of train loss from Figure 1(b) and 2(b). When using with-replacement sampling, DSGD achieves higher test accuracy than that of DSGD-M and DSignSGD, but exhibits the worst training-loss-performance from Figure 3- 4.
Furthermore, in all the compared methods, the consensus error diminishes to 0 throughout the iterations. These results illustrate that our framework is able to yield efficient stochastic subgradient-based methods for decentralized training tasks.
@@ 5.3 Performance comparison between DSGD and DSGD-T
The numerical results on the CIFAR-100 and CIFAR-10 datasets are presented in Figures 5- 8, respectively. It can be concluded that DSGD-T demonstrates slightly better performance than DSGD in all aspects under all sampling strategies, especially in train loss. Therefore, we can conclude that DSGD-T achieves high efficiency in decentralized training tasks while enjoying convergence guarantees from our framework (DSM).
@@
Figure 5: Numerical performance comparison of DSGD and DSGD-T in training ResNet50 on CIFAR-100 dataset using random reshuffling strategy.
@@
Figure 6: Numerical performance comparison of DSGD and DSGD-T in training ResNet50 on CIFAR-10 dataset using random reshuffling strategy.
(a) Test Accuracy
(b) Train Loss
(c) Consensus Error
Figure 7: Numerical performance comparison of DSGD and DSGD-T in training ResNet50 on CIFAR-100 dataset using with-replacement sampling strategy.
(a) Test Accuracy
(b) Train Loss
(c) Consensus Error
Figure 8: Numerical performance comparison of DSGD and DSGD-T in training ResNet50 on CIFAR-10 dataset using with-replacement sampling strategy.
@@ 6 Conclusion
Our work sets several directions for future research in the field of nonsmooth nonconvex decentralized optimization. Existing ODE approaches are limited to establishing asymptotic convergence properties or the convergence rate for the continuous-time trajectories of the corresponding differential inclusion [NEURIPS2022_a9077da4, bolte2025inexact]. Consequently, deriving explicit convergence rates for various nonsmooth decentralized stochastic subgradient methods (such as DSGD, DSGT, and their momentum variants) remains a significant challenge. While our analysis establishes the convergence of DSGT in nonsmooth nonconvex settings, it reveals that DSGT and DSGD are corresponding to the same limiting differential inclusion. Empirically, however, Figures 5-8 demonstrate that DSGT consistently converges faster than DSGD. Bridging this gap between theory and practice, especially in establishing explicit convergence rates that distinguish between DSGT and DSGD in the nonsmooth nonconvex regime, remains a critical direction for future research.
Further attention is also needed on the convergence of decentralized methods with time-varying networks, as well as those that incorporate asynchronous or compressed communication. Additionally, we are also eager to explore routes that presume relaxed conditions for convergence analysis of decentralized methods in solving (DOP). For instance, closed-measure approaches [bianchi2021closed, bolte2022long, le2024nonsmooth] are developed for establishing the convergence properties of stochastic subgradient-based methods, with relaxed conditions on the step-sizes and evaluation noises. How to employ these closed-measure approaches for analyzing decentralized stochastic subgradient-based methods is worthy of investigating.
@@ Acknowledgement
$\mathcal{U}_{k}({\bm{x}})$ $=\left(D_{F_{1},\mathcal{B}_{1,k+1}}({\bm{x}}_{(1)}),\ldots,D_{F_{d},\mathcal{B}_{d,k+1}}({\bm{x}}_{(d)})\right)^{\top}$
Note that $(F_{i}(\cdot,{\bm{s}}_{l}),D_{F_{i}}(\cdot,{\bm{s}}_{l}))$ has a $\mathcal{C}^{r}$ variational stratification [bolte2021conservative, Theorem 4], hence the mappings $\{\mathcal{U}_{k}\}$ are almost everywhere $\mathcal{C}^{1}$ over $\mathbb{R}^{nd}$, denoted by $\mathcal{U}_{k}({\bm{x}})=\{u_{k}({\bm{x}})\},$ for a.e. ${\bm{x}}\in\mathbb{R}^{nd}$. Furthermore, let
$\mathcal{Y}$ $=\left\{{\bm{x}}\in\mathbb{R}^{nd}:D_{F_{i}}({\bm{x}};{\bm{s}}_{l})\neq\{\nabla F_{i}({\bm{x}},{\bm{s}}_{l})\},\text{ for some }i\in[N]\text{ and some }{\bm{s}}_{l}\in\mathcal{S}_{i}\right\},$
@@
is an open covering of $\{{\bm{x}}\in\Gamma_{k}:\det({\bm{J}}_{c,k}({\bm{x}}))\neq 0\}$. Based on Lindelof’s lemma [kelley2017general], there exists countable $\{{\bm{x}}^{i}\}_{i\in\mathbb{N}_{+}}\subseteq\{{\bm{x}}\in\Gamma_{k}:\det({\bm{J}}_{c,k}({\bm{x}}))\neq 0\}$, such that
@@
Let $\mathcal{S}:=\cap_{k\geq p}\mathcal{S}_{k}$, $A_{0}:=A=\mathcal{Y},A_{k+1}:=\{{\bm{x}}\in\mathbb{R}^{nd}:{\bm{P}}{\bm{z}}-cq_{k}({\bm{z}})\in A\}$, and $\hat{A}:=\cup_{k\geq 0}A_{k}$. It can be inferred that $\mathcal{S}$ and $\hat{A}^{c}$ are full-measure subsets in $\mathbb{R}$ and $\mathbb{R}^{nd}$, respectively. For any $c\in\mathcal{S}$, and any ${\bm{x}}\in\hat{A}^{c}$, we have ${\bm{P}}{\bm{x}}-cq_{k}({\bm{x}})\in A^{c}$. Employing Fubini’s theorem, the set $\{c\in\mathbb{R},{\bm{x}}\in\mathbb{R}^{nd},\cup_{k\geq 0}\{{\bm{P}}{\bm{x}}-cq_{k}({\bm{x}})\}\subseteq A^{c}\}$ is full-measure in $\mathbb{R}\times\mathbb{R}^{nd}$, and thus, there exists a full-measure subset $\hat{\mathcal{S}}\times\hat{\mathcal{K}}$, whenever $(c,{\bm{x}}_{0})\in\hat{\mathcal{S}}\times\hat{\mathcal{K}}$, it follows that $\cup_{k\geq 0}\{{\bm{P}}{\bm{x}}-cq_{k}({\bm{x}})\}\subseteq A^{c}$, i.e. $\{{\bm{x}}_{k}\}\in\mathcal{Y}^{c}$.
@@
$\left\{\begin{aligned} {\bm{g}}_{k}={}&\left(\nabla F_{1,\mathcal{B}_{1,k+1}}({\bm{x}}_{1,k}),\ldots,\nabla F_{d,\mathcal{B}_{d,k+1}}({\bm{x}}_{d,k})\right)^{\top},\\ {\bm{x}}_{k+1}&\in{}({\bm{W}}\otimes{\bm{I}}_{n}){\bm{x}}_{k}-c\hat{\eta}_{k}{\bm{g}}_{k}\end{aligned}\right$.
@@
$\mathcal{U}_{1,k}({\bm{x}},{\bm{y}})$ $=\left(\partial\hat{\phi}({\bm{y}}),{\bm{0}}\right)$
$\mathcal{U}_{2,k}({\bm{x}},{\bm{y}})$ $=\left({\bm{0}},\tau({\bm{W}}\otimes{\bm{I}}_{n}){\bm{y}}-\tau\left(D_{F_{1},\mathcal{B}_{1,k+1}}({\bm{x}}_{(1)}),\ldots,D_{F_{d},\mathcal{B}_{d,k+1}}({\bm{x}}_{(d)})\right)^{\top}\right)$
where $\hat{\phi}({\bm{y}}):=(\phi({\bm{y}}_{(1)}),\ldots,\phi({\bm{y}}_{(d)}))^{\top}$, $\phi:\mathbb{R}^{n}\to\mathbb{R}$ is convex and admits a unique minimizer at $\bm{0}$}. Notice that $(F_{i}(\cdot,{\bm{s}}_{l}),D_{F_{i}(\cdot,{\bm{s}}_{l})})$ has a $\mathcal{C}^{r}$ variational stratification [bolte2021conservative, Theorem 4], and $\hat{\phi}$ is differentiable on a full-measure subset of $\mathbb{R}^{nd}$, thus the mappings $\{\mathcal{U}_{j,k}\},j=1,2$ are almost everywhere $\mathcal{C}^{1}$ over $\mathbb{R}^{2nd}$, the corresponding almost-everywhere-$\mathcal{C}^{1}$ mappings are denoted by $u_{j,k}({\bm{z}}),j=1,2$. Furthermore, let
$\mathcal{Y}$ $=\left\{({\bm{x}},{\bm{y}})\in\mathbb{R}^{2nd}:\hat{\phi}({\bm{y}})\text{ is not differentiable },D_{F_{i,l}}({\bm{x}})\neq\{\nabla F_{i}({\bm{x}},{\bm{s}}_{l})\},\text{ for some }i\in[N]\right$.
 $\left.\text{ and some }{\bm{s}}_{l}\in\mathcal{S}_{i}\right\},$
@@
$\left\{\begin{aligned} {\bm{x}}_{k+1}&\in{}({\bm{W}}\otimes{\bm{I}}_{n}){\bm{x}}_{k}-c\hat{\eta}_{k}\nabla\phi({\bm{y}}_{k}),\\ {\bm{g}}_{k+1}={}&\left(\nabla F_{1,\mathcal{B}_{1,k+1}}({\bm{x}}_{1,k+1}),\ldots,\nabla F_{d,\mathcal{B}_{d,k+1}}({\bm{x}}_{d,k+1})\right)^{\top},\\ {\bm{y}}_{k+1}&\in{}(1-c\tau\hat{\eta}_{k})({\bm{W}}\otimes{\bm{I}}_{n}){\bm{y}}_{k}+c\tau\hat{\eta}_{k}{\bm{g}}_{k+1}.\end{aligned}\right$.
Appendix B Synthetic numerical examples on avoiding spurious critical points
In this part, we present numerical experiments on a synthetic example to illustrate that, under random initialization, Algorithm 2 converges to a Clarke-critical point with probability $1$.
We consider a four-node ring network with a non-singular mixing matrix
$W=\frac{1}{3}\begin{pmatrix}1&1&0&1\\ 1&1&1&0\\ 0&1&1&1\\ 1&0&1&1\end{pmatrix}$.
Moreover, the local objective functions are defined as
$f_{1}(x,y)$ $=\mathrm{ReLU}(x)+\mathrm{ReLU}(y),$
$f_{2}(x,y)$ $=\mathrm{ReLU}(x)+\mathrm{ReLU}(y),$
$f_{3}(x,y)$ $=\mathrm{ReLU}(-x)+\mathrm{ReLU}(-y),$
$f_{4}(x,y)$ $=\mathrm{ReLU}(-x-1)+(x+1)-\mathrm{ReLU}(x+1)+\mathrm{ReLU}(-x-2)+(x+2)-\mathrm{ReLU}(x+2)$
$\quad+\mathrm{ReLU}(-y-1)+(y+1)-\mathrm{ReLU}(y+1),$
respectively.
Note that $f_{4}(x,y)$ is identically zero. However, by following the same theoretical analysis in [NEURIPS2022_a9077da4], automatic differentiation (AD) yields $\partial_{x}f_{4}(-1,y)=1$, $\partial_{x}f_{4}(-2,y)=1$, and $\partial_{y}f_{4}(x,-1)=1$ for any $x\in\mathbb{R}$. As a result, AD introduces spurious stationary points for $f=f_{1}+f_{2}+f_{3}+f_{4}$, which are
(1,0),(0,1),(2,0),(1,1),(2,1).(-1,0),\ (0,-1),\ (-2,0),\ (-1,-1),\ (-2,-1).
In our experiments, we randomly initialize $(x_{0},y_{0})\sim\mathcal{N}\!\bigl((0,0),I_{2}\bigr)$ in Algorithm 2, and in each epoch we compute only one full-batch local gradient using AD. Figure 9 reports a randomly sampled iterate trajectory and the percentages over $100$ independent runs that converge to different critical points. A run is declared to converge to a critical point if the squared distance from its final iterate to that point is below $10^{-4}$. As shown in Figure 9, almost all runs generated by Algorithm 2 converge to the true Clarke-critical point $(0,0)$, which is consistent with Theorem 4.8.
(a) Iterate trajectory
(b) The percentages on different types of the final results
Figure 9: Numerical results of Algorithm 2 on a toy example.
Appendix C Developing Decentralized Generalized ADAM Methods with Convergence Guarantees
Adam [adam] uses a momentum term and an exponential moving average of squared gradients as an adaptive scaling factor, thereby applying coordinate-wise adaptive step sizes and updating parameters with bias correction. After this work, a number of efficient variants are developed, such as AdaBelief [zhuang2020adabelief], AMSGrad [reddi2019convergence], decentralized AMSGrad [chen2023convergence], etc. In the single-agent setting, [xiao2023adam, ding2023adam] investigate the global convergence properties of Adam-family methods for nonsmooth objectives via a delicately constructed Lyapunov function.
In this part, we present a decentralized generalized Adam method in Algorithm 4, which fits into (DSM). Notice $b_{i,1},\ldots,b_{i,s}$ in Line 9 represent the block sizes in Adam-mini at node $i$.
Let ${\bm{M}}_{k}:=[{\bm{m}}_{1,k},\ldots,{\bm{m}}_{d,k}]$, ${\bm{V}}_{k}:=[{\bm{v}}_{1,k},\ldots,{\bm{v}}_{d,k}]$. We have the following theorem illustrating the convergence properties of Algorithm 4.
Theorem C.1.
Suppose Assumption 4.1 holds. $\{{\bm{X}}_{k}\}$ is generated by Algorithm 4.
  1. (1)
  2. (2)
    and
Algorithm 4 DGAdam for solving (DOP).
1:Initial point ${\bm{x}}_{0}\in\mathbb{R}^{n}$, initial momentum terms ${\bm{m}}_{0},{\bm{v}}_{0}\in\mathbb{R}^{n}$, momentum parameters $\tau_{1},\tau_{2}>0$, an auxiliary function $\phi:\mathbb{R}^{n}\to\mathbb{R}$, and a mixing matrix $\bm{W}$}.
2:for all $i\in[d]$ in parallel do
3:  Set $k\leftarrow 0$. Initialize ${\bm{x}}_{i,k}={\bm{x}}_{0},{\bm{m}}_{i,k}={\bm{m}}_{0},$;
4:  while not terminated do
5:   Choose the step-size $\eta_{k}$;
6:   Communicate and update the local variable
${\bm{x}}_{i,k+1}\in\sum_{j\in\mathcal{N}_{i}}{\bm{W}}(i,j){\bm{x}}_{j,k}-\eta_{k}\frac{1}{\sqrt{{\bm{v}}_{i,k}+\epsilon}}\odot({\bm{m}}_{i,k}+\rho{\bm{d}}_{i,k});$
7:   Compute ${\bm{d}}_{i,k+1}\in\frac{1}{|\mathcal{B}_{i,k+1}|}\sum_{{\bm{s}}_{l}\in\mathcal{B}_{i,k+1}}D_{F_{i,l}}({\bm{x}}_{i,k+1})$;
8:   Randomly select a mini-batch $\mathcal{B}_{i,k}\subseteq\mathcal{S}_{i}$;
9:   Communicate and update momentum
${\bm{m}}_{i,k+1}=(1-\tau_{1}\eta_{k})\sum_{j\in\mathcal{N}_{i}}{\bm{W}}(i,j){\bm{m}}_{j,k}+\tau_{1}\eta_{k}{\bm{d}}_{i,k+1};$
10:   Communicate and update second-moment estimate by one of the formats
${\bm{v}}_{i,k+1}$ $\in(1-\tau_{2}\eta_{k})\sum_{j\in\mathcal{N}_{i}}{\bm{W}}(i,j){\bm{v}}_{j,k}+\tau_{2}\eta_{k}{\bm{d}}_{i,k+1}\odot{\bm{d}}_{i,k+1};\hskip 90.00014pt\triangleleft\text{ Adam }$
${\bm{v}}_{i,k+1}$ $\in(1-\tau_{2}\eta_{k})\sum_{j\in\mathcal{N}_{i}}{\bm{W}}(i,j){\bm{v}}_{j,k}+\tau_{2}\eta_{k}\mathrm{Diag}(\frac{{\bm{1}}{\bm{1}}^{\top}}{b_{i,1}},\ldots,\frac{{\bm{1}}{\bm{1}}^{\top}}{b_{i,s}}){\bm{d}}_{i,k+1}\odot{\bm{d}}_{i,k+1};\triangleleft\text{ Adam-mini }$
11:   Set $k\leftarrow k+1$;
12:  end while
13:end for
14:return ${\bm{X}}_{k}:=[{\bm{x}}_{1,k},\ldots,{\bm{x}}_{d,k}]$;
Proof.
The compact format of Algorithm 4 with Adam update can be written as
${\bm{X}}_{k+1}\in{}$ ${\bm{X}}_{k}{\bm{W}}-\eta_{k}\frac{1}{\sqrt{{\bm{V}}_{k}+\epsilon{\bm{I}}}}\odot({\bm{M}}_{k}+\rho{\bm{D}}_{k}),$ (C.1)
${\bm{D}}_{k+1}={}$ $[{\bm{d}}_{1,k+1},\ldots,{\bm{d}}_{d,k+1}],$
${\bm{M}}_{k+1}={}$ $(1-\tau_{1}\eta_{k}){\bm{M}}_{k}{\bm{W}}+\tau_{1}\eta_{k}{\bm{D}}_{k+1},$
${\bm{V}}_{k+1}\in{}$ $(1-\tau_{2}\eta_{k}){\bm{V}}_{k}{\bm{W}}+\tau_{2}\eta_{k}{\bm{D}}_{k+1}\odot{\bm{D}}_{k+1},$
However, if we directly define ${\bm{Z}}_{k}:=\left[\begin{smallmatrix}{\bm{X}}_{k}\\ {\bm{M}}_{k}\\ {\bm{V}}_{k}\end{smallmatrix}\right]$, we cannot find a function $\Phi_{i}$ that satisfies Assumption 3.1. To address this issue, we introduce an auxiliary update scheme to Algorithm 4 parameterized by $K>0$.
${\bm{X}}_{k+1}\in{}$ ${\bm{X}}_{k}{\bm{W}}-\eta_{k}\frac{1}{\sqrt{{\bm{V}}_{k}+\epsilon{\bm{I}}}}\odot({\bm{M}}_{k}+\rho{\bm{D}}_{k}),$ (C.2)
${\bm{D}}_{k+1}={}$ $[{\bm{d}}_{1,k+1},\ldots,{\bm{d}}_{d,k+1}],$
${\bm{M}}_{k+1}={}$ $(1-\tau_{1}\eta_{k}){\bm{M}}_{k}{\bm{W}}+\tau_{1}\eta_{k}{\bm{D}}_{k+1},$
$\frac{1}{d}{\bm{V}}_{k+1}{\bm{1}}_{d}{\bm{1}}^{\top}_{d}\in{}$ $(1-\tau_{2}\eta_{k})\frac{1}{d}{\bm{V}}_{k+1}{\bm{1}}_{d}{\bm{1}}^{\top}_{d}{\bm{W}}+\tau_{2}\eta_{k}\frac{1}{d}{\bm{D}}_{k+1}\odot{\bm{D}}_{k+1}{\bm{1}}_{d}{\bm{1}}^{\top}_{d}\mathbb{1}_{K}({\bm{V}}_{k}),$
where $\mathbb{1}_{K}$ denotes the indication function of the subset $\{{\bm{V}}:\|{\bm{V}}\|\leq K\}$. Denote the filtration $\mathcal{F}_{k}:=\sigma(\{{\bm{X}}_{l}:l\leq k\})$, $D_{F_{\mathcal{B}_{i,k+1}}}({\bm{x}}_{i,k+1}):=\frac{1}{|\mathcal{B}_{i,k+1}|}\sum_{{\bm{s}}_{l}\in\mathcal{B}_{i,k+1}}D_{F_{i,l}}({\bm{x}}_{i,k+1})$, and let